Shift report 20010829 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Aug 29 2001 - 16:06:01 EDT

  • Next message: Flemming Videbaek: "today's access"

    ------------------------------------------------------------
    Date: Wed Aug 29 16:06:00 2001, Shift: 08:00-16:00
    Supervisor: Alexandru JIPA, with: -
    Log book#8 pages 37 - 38
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beams up to 12.30. After this hour - maintenance.
    2. After discussions with Flemming Videbaek on detector status
    - one run (4495), without magnetic filed and without TFCAL; T3
     - second module OFF.
    3. Works at T3 and HV for different detectors. T3 was removed
     from experimental set-up. 
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Wed Aug 29 2001 - 16:07:45 EDT